Ques) A capacitor of capacitance  C1 = 1uF charged upto a voltage of V=110V is connected in parallel to the terminals of  a circuit consisting of a two uncharged capapcitors connected in series and possessing the capapcitance  C2 = 2 uF and C3 = 3uF. What charge will flow through the connecting wires??

Grade:12

1 Answers

AskiitianExpert Shine
14 years ago

Hi

After connecting all three capacitors become in parallel.

      so before connecting Total charge=110 uC =Total charge on all 3 capacitors after connecting =q1+q2+q3

      also after connection, voltage across all three is equal so
                  q1=q2/2=q3/3

         

Using these equations we get

           q1=18.3333 uC

           So, charge flown through connecting wires= 110uC -q1.

hence the answer is 91.667 uC.